3 years ago, Ade was 5 years old and his mother is 5 times his age.
How many year' will his mother be in 4 years?
How many years' will Ade be in 4 years?

Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)


3 years ago, Ade was 5 years old and his mother is 5 times his age.
A-3 = 5M+5
-----
A = 5M + 8
M = (A-8)/5
----------------
How many years will his mother be in 4 years?
M+4 = (1/5)A - (8/5) + 4 = (1/5)A + 12/5
-------------------------------------------------
How many years' will Ade be in 4 years?
A + 4 = 5M+8
------
Cheers,
Stan H.
----------------


Answer by MathTherapy(10552)

3 years ago, Ade was 5 years old and his mother is 5 times his age.
How many year' will his mother be in 4 years?
How many years' will Ade be in 4 years?
If 3 years ago he was 5, then how old is he now?
If his mother is 5 times his age, how old is the mother?
Is this so hard?
Now you should be able to answer the questions you posed!